OpenStudy (anonymous):

help please dont know how to do this. solve the equation, check for extraneous solutions. square root 2x+9- square root x =3

OpenStudy (anonymous):

\[\sqrt{2x+9}-\sqrt{x}=3\]

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

can anyone help me please